# Intuitionistic Fuzzy Universum Twin Support Vector Machine for Imbalanced Data

### v1

@@ Abstract
One of the major difficulties in machine learning methods is categorizing datasets that are imbalanced. This problem may lead to biased models, where the training process is dominated by the majority class, resulting in inadequate representation of the minority class. Universum twin support vector machine (UTSVM) produces a biased model towards the majority class, as a result, its performance on the minority class is often poor as it might be mistakenly classified as noise. Moreover, UTSVM is not proficient in handling datasets that contain outliers and noises. Inspired by the concept of incorporating prior information about the data and employing an intuitionistic fuzzy membership scheme, we propose intuitionistic fuzzy universum twin support vector machines for imbalanced data (IFUTSVM-ID). We use an intuitionistic fuzzy membership scheme to mitigate the impact of noise and outliers. Moreover, to tackle the problem of imbalanced class distribution, data oversampling and undersampling methods are utilized. Prior knowledge about the data is provided by universum data. This leads to better generalization performance. UTSVM is susceptible to overfitting risks due to the omission of the structural risk minimization (SRM) principle in their primal formulations. However, the proposed IFUTSVM-ID model incorporates the SRM principle through the incorporation of regularization terms, effectively addressing the issue of overfitting. We conduct a comprehensive evaluation of the proposed IFUTSVM-ID model on benchmark datasets from KEEL and compare it with existing baseline models. Furthermore, to assess the effectiveness of the proposed IFUTSVM-ID model in diagnosing Alzheimer’s disease (AD), we applied them to the Alzheimer’s Disease Neuroimaging Initiative (ADNI) dataset. Experimental results showcase the superiority of the proposed IFUTSVM-ID models compared to the baseline models.
keywords:
@@
[…]
[inst1]organization=Department of Mathematics,addressline=Indian Institute of Technology Indore, city=Simrol, Indore, postcode=453552, country=India
1 Introduction
@@
Support vector machines (SVMs) [5] are advanced machine learning models that maximize the margin between two classes in a classification problem, aiming to find the optimal hyperplane between two parallel supporting hyperplanes. SVM has proven its remarkable utility across diverse domains such as fault detection [47], biomedicine [23], text categorization [14], and disease diagnosis [40, 30] and so on. SVM integrates the structural risk minimization (SRM) principle within its optimization framework, thereby enhancing its generalization capabilities by minimizing an upper bound of the generalization error. SVM solves one large quadratic programming problem (QPP), resulting in escalated computational complexity, which renders it less suitable for large-scale datasets. To enhance the efficiency of SVM, Jayadeva et al. [13] proposed twin support vector machine (TSVM). TSVM generates a pair of non-parallel hyperplanes, with each hyperplane deliberately situated in close proximity to the data points belonging to one class while ensuring a minimum separation distance of at least one unit from the data points of the other class. TSVM solves two smaller-sized QPPs instead of a single large QPP, making TSVM four times faster than the standard SVM. However, TSVM faced two notable challenges: the requirement for matrix inversions and the absence of the SRM principle in its formulation, which posed significant obstacles to its effectiveness. Shao et al. [35] introduced twin-bounded SVM (TBSVM) which incorporated a regularization term in their formulation, allowing the principle of SRM to be employed. In recent years, numerous variants of TSVM have been proposed, including granular ball TSVM [26, 28], robust energy-based LTSVM (RELTSVM) [36], multiview learning with twin parametric margin SVM (MvTPMSVM) [25] and others.
@@
Several variants of the SVM model have been developed with the aim of improving its generalization performance; however, specific attributes of the original data may have been disregarded. In practical applications, it is common to encounter data that do not pertain to our target class but are within the same domain. These data, known as Universum data, can be utilized to enhance the learning process. Universum data consists of samples that do not belong to any of the relevant classes, providing insights into the distribution of the data. By integrating universum data into the SVM classifier, Weston et al. [41] proposed universum SVM (USVM). In USVM, the universum data points are integrated within the $\epsilon$-insensitive tube positioned between the binary classes. As a result, the resulting model demonstrates enhanced generalization performance. Universum twin support vector machine (UTSVM) [24] has been introduced to decrease the training time of USVM. Furthermore, many variants of SVM and TSVM based on universum data have been proposed over the years to enhance the efficiency of the models such as universum least square TSVM [42], non-parallel hyperplane universum SVM (U-NHSVM) [48], lagrangian-based approach for universum twin-bounded SVM [21], multi-view UTSVM with Insensitive Pinball Loss (Pin-MvUTSVM) [20], universum twin support vector machine with pinball loss function (Pin-UTSVM) [9], $p$-norm least-squares Universum twin bounded support vector machine (LSp-UTBSVM) [22], twin parametric margin SVM based on Universum data (UTPMSVM) [11] and so on.
@@
While TSVM and its variants effectively handle the computational complexity challenges posed by SVM, both SVM and TSVM encounter difficulties when dealing with class-imbalanced data, particularly in datasets containing noise and outliers. The accuracy of SVM and TSVM models’ predictions is significantly affected by noisy data, outliers, and imbalanced datasets. The issue of class imbalance occurs when there is a notable difference in the number of samples belonging to one particular class compared to the others within a dataset. Although a uniform weighting scheme is applied to each data sample during the generation of the optimal classifier by SVM and TSVM, they are still susceptible to issues such as noise, outliers, and class imbalance. The fuzzy theory has proven to be effective in mitigating the adverse impacts of noise or outliers. Many fuzzy scheme-based approaches in SVM have been introduced so far for the classification of noisy data points [18, 2, 39]. In classification problems where there is a significant imbalance between samples of majority and minority class, SVM classification models often exhibit a biasness towards the majority class samples, which can lead to incorrect classification of samples of the minority class. To address the issue of class imbalance, an entropy-based approach proposed in [8]. This approach assigns higher membership to each individual data point lying on the supporting hyperplane. Richhariya and Tanveer [32] employed a fuzzy membership scheme in both FUSVM and FUTSVM considering the probability of uncertainty of data points belonging or not belonging to the respective class. In the presence of noise and outliers, classification models encounter difficulties with imbalanced classes. In situations such as disease diagnosis [30, 40], the primary objective is to accurately classify minority class samples. Various models are available to tackle the problem of imbalanced classes, such as fuzzy SVM (FSVM) for class imbalance learning [3, 17], which assigns fuzzy membership weights to the data samples. Further techniques have been developed to handle class imbalance, including random undersampling approach [19, 45] and synthetic minority oversampling technique (SMOTE) [4]. To address class imbalance learning, robust fuzzy LSTSVM (RFLSTSVM-CIL) [31] utilized the imbalance ratio to the data points to generate fuzzy weights, leading to an improvement in generalization performance. In order to mitigate the impact of noise, some approaches have been proposed for class imbalance problems in [43, 44]. TSVM remains sensitive to datasets that contain noise and outliers. In [1], a variant of the fuzzy set known as the intuitionistic fuzzy set (IFS) is proposed. IFS incorporates both membership and non-membership functions, whereas a fuzzy set is only based on membership functions. This is the primary difference between IFS and a fuzzy set. In order to reduce the impact of noise and outliers, a combination of TSVM and IFS was proposed in [29].
@@
The remainder structure of this paper as follows: Section 2 briefly explains the formulation of UTSVM and intuitionistic fuzzy membership scheme. In Section 3, we derive the mathematical formulation of the proposed IFUTSVM-ID model. The time complexity of the proposed IFUTSVM-ID model is detailed in Section 4. Numerical experiment comparisons are demonstrated in Section 5. Conclusions and future directions are presented in section 6.
2 Related works
@@
In this section, we briefly discuss the intuitionistic fuzzy membership Scheme and the mathematical formulation of UTSVM.
2.1 Intuitionistic Fuzzy Membership Scheme

### v2

# Robust Universum Twin Support Vector Machine for Imbalanced Data

@@ Abstract
One of the major difficulties in machine learning methods is categorizing datasets that are imbalanced. This problem may lead to biased models, where the training process is dominated by the majority class, resulting in inadequate representation of the minority class. Universum twin support vector machine (UTSVM) produces a biased model towards the majority class, as a result, its performance on the minority class is often poor as it might be mistakenly classified as noise. Moreover, UTSVM is not proficient in handling datasets that contain outliers and noises. Inspired by the concept of incorporating prior information about the data and employing an intuitionistic fuzzy membership scheme, we propose intuitionistic fuzzy universum twin support vector machines for imbalanced data (IFUTSVM-ID) by enhancing overall robustness. We use an intuitionistic fuzzy membership scheme to mitigate the impact of noise and outliers. Moreover, to tackle the problem of imbalanced class distribution, data oversampling and undersampling methods are utilized. Prior knowledge about the data is provided by universum data. This leads to better generalization performance. UTSVM is susceptible to overfitting risks due to the omission of the structural risk minimization (SRM) principle in their primal formulations. However, the proposed IFUTSVM-ID model incorporates the SRM principle through the incorporation of regularization terms, effectively addressing the issue of overfitting. We conduct a comprehensive evaluation of the proposed IFUTSVM-ID model on benchmark datasets from KEEL and compare it with existing baseline models. Furthermore, to assess the effectiveness of the proposed IFUTSVM-ID model in diagnosing Alzheimer’s disease (AD), we applied them to the Alzheimer’s Disease Neuroimaging Initiative (ADNI) dataset. Experimental results showcase the superiority of the proposed IFUTSVM-ID models compared to the baseline models. The supplementary material of the paper can be accessed using the following link: https://github.com/AnonymousAuthor0011/IFUTSVM-ID.git.
Index Terms:
@@
I Introduction
@@
Support vector machines (SVMs) [1] are advanced machine learning models that maximize the margin between two classes in a classification problem, aiming to find the optimal hyperplane between two parallel supporting hyperplanes. SVM has proven its remarkable utility across diverse domains such as fault detection [2], text categorization [3], and disease diagnosis [4, 5] and so on. SVM integrates the structural risk minimization (SRM) principle within its optimization framework, thereby enhancing its generalization capabilities by minimizing an upper bound of the generalization error. SVM solves one large quadratic programming problem (QPP), resulting in escalated computational complexity, which renders it less suitable for large-scale datasets. To enhance the efficiency of SVM, Jayadeva et al. [6] proposed twin support vector machine (TSVM). TSVM generates a pair of non-parallel hyperplanes, with each hyperplane deliberately situated in close proximity to the data points belonging to one class while ensuring a minimum separation distance of at least one unit from the data points of the other class. TSVM solves two smaller-sized QPPs instead of a single large QPP, making TSVM four times faster than the standard SVM. However, TSVM faced two notable challenges: the requirement for matrix inversions and the absence of the SRM principle in its formulation, which posed significant obstacles to its effectiveness. Shao et al. [7] introduced twin-bounded SVM (TBSVM), which incorporated a regularization term in their formulation, allowing the principle of SRM to be employed. In recent years, numerous variants of TSVM have been proposed, including granular ball TSVM [8, 9], robust energy-based LTSVM (RELTSVM) [10], multiview learning with twin parametric margin SVM (MvTPMSVM) [11], Wave-MvSVM [12], twin restricted kernel machine (TRKM) [13], one class restricted kernel machine (OCRKM) [14] and others.
@@
Several variants of the SVM model have been developed with the aim of improving its generalization performance; however, specific attributes of the original data may have been disregarded. In practical applications, it is common to encounter data that do not pertain to our target class but are within the same domain. These data, known as Universum data, can be utilized to enhance the learning process. Universum data consists of samples that do not belong to any of the relevant classes, providing insights into the distribution of the data. By integrating universum data into the SVM classifier, Weston et al. [15] proposed universum SVM (USVM). In USVM, the universum data points are integrated within the $\epsilon$-insensitive tube positioned between the binary classes. As a result, the resulting model demonstrates enhanced generalization performance. Universum twin support vector machine (UTSVM) [16] has been introduced to decrease the training time of USVM. Furthermore, many variants of SVM and TSVM based on universum data have been proposed over the years to enhance the efficiency of the models such as non-parallel hyperplane universum SVM (U-NHSVM) [17], lagrangian-based approach for universum twin-bounded SVM [18], multi-view UTSVM with Insensitive Pinball Loss (Pin-MvUTSVM) [19], twin parametric margin SVM based on Universum data (UTPMSVM) [20] and so on.
@@
While TSVM and its variants effectively handle the computational complexity challenges posed by SVM, both SVM and TSVM encounter difficulties when dealing with class-imbalanced data, particularly in datasets containing noise and outliers. The accuracy of SVM and TSVM models’ predictions is significantly affected by noisy data, outliers, and imbalanced datasets. The issue of class imbalance occurs when there is a notable difference in the number of samples belonging to one particular class compared to the others within a dataset. Although a uniform weighting scheme is applied to each data sample during the generation of the optimal classifier by SVM and TSVM, they are still susceptible to issues such as noise, outliers, and class imbalance. The fuzzy theory has proven to be effective in mitigating the adverse impacts of noise or outliers. Many fuzzy scheme-based approaches in SVM have been introduced so far for the classification of noisy data points [21, 22]. In classification problems where there is a significant imbalance between samples of majority and minority class, SVM classification models often exhibit a biasness towards the majority class samples, which can lead to incorrect classification of samples of the minority class. To address the issue of class imbalance, an entropy-based approach proposed in [23]. This approach assigns higher membership to each individual data point lying on the supporting hyperplane. Richhariya and Tanveer [24] employed a fuzzy membership scheme in both FUSVM and FUTSVM considering the probability of uncertainty of data points belonging or not belonging to the respective class. In the presence of noise and outliers, classification models encounter difficulties with imbalanced classes. In situations such as disease diagnosis [5], the primary objective is to accurately classify minority class samples. Various models are available to tackle the problem of imbalanced classes, such as fuzzy SVM (FSVM) for class imbalance learning [25], which assigns fuzzy membership weights to the data samples. To address class imbalance learning, robust fuzzy LSTSVM (RFLSTSVM-CIL) [26] utilized the imbalance ratio to the data points to generate fuzzy weights, leading to an improvement in generalization performance. In order to mitigate the impact of noise, some approaches have been proposed for class imbalance problems in [27]. TSVM remains sensitive to datasets that contain noise and outliers. In [28], a variant of the fuzzy set known as the intuitionistic fuzzy set (IFS) is proposed. IFS incorporates both membership and non-membership functions, whereas a fuzzy set is only based on membership functions. This is the primary difference between IFS and a fuzzy set. In order to reduce the impact of noise and outliers, a combination of TSVM and IFS was proposed in [29].
@@
The remainder structure of this paper as follows: Section II briefly explains the formulation of UTSVM and intuitionistic fuzzy membership scheme. In Section III, we derive the mathematical formulation of the proposed IFUTSVM-ID model. Numerical experiment comparisons are demonstrated in Section IV. Conclusions and future directions are presented in section V.
II Related works
@@
In this section, we go through the mathematical formulation of UTSVM. The Intuitionistic fuzzy membership Scheme is discussed in Section S.I of the Supplementary Material.
II-A Universum Twin Support Vector Machine (UTSVM)
$\underset{w_{1},b_{1},\xi_{1},\psi_{1}}{min}\hskip 14.22636pt\leavevmode\nobreak$ $\frac{1}{2}\|X_{1}w_{1}+e_{1}b_{1}\|^{2}+c_{1}e_{2}^{t}\xi_{1}+c_{u}e_{u}^{t}\psi_{1}$
s.t. $-\left(X_{2}w_{1}+e_{2}b_{1}\right)\geq e_{2}-\xi_{1},$
$\left(Uw_{1}+e_{u}b_{1}\right)\geq(-1+\epsilon)e_{u}-\psi_{1},$
$\hskip 17.07182pt\xi_{1}\geq 0,\psi_{1}\geq 0,$ (1)
and
$\underset{w_{2},b_{2},\xi_{2},\psi_{2}}{min}\hskip 14.22636pt\leavevmode\nobreak$ $\frac{1}{2}\|X_{2}w_{2}+e_{2}b_{2}\|^{2}+c_{2}e_{1}^{t}\xi_{2}+c_{u}e_{u}^{t}\psi_{2}$
s.t. $\left(X_{1}w_{2}+e_{1}b_{2}\right)\geq e_{1}-\xi_{2},$
$-\left(Uw_{2}+e_{u}b_{2}\right)\geq(-1+\epsilon)e_{u}-\psi_{2},$
$\hskip 17.07182pt\xi_{2}\geq 0,\psi_{2}\geq 0,$ (2)
here, $c_{i},(i=1,2,u)$ are positive parameters, $\xi_{1},\xi_{2},\psi_{1},\psi_{2}$ are slacks variables. The dual formulations of problems (II-A) and (II-A) can be obtaind as:
$\underset{\alpha_{1},\mu_{1}}{max}\hskip 14.22636pt\leavevmode\nobreak$ $e_{2}^{t}\alpha_{1}-\frac{1}{2}(\alpha_{1}^{t}E-\mu_{1}^{t}F)(GG^{t})^{-1}(E^{t}\alpha_{1}-F^{t}\mu_{1})$
$+(\epsilon-1)e_{u}^{t}\mu_{1}$
s.t $0\leq\mu_{1}\leq c_{u},$
$0\leq\alpha_{1}\leq c_{1},$
and
$\underset{\alpha_{2},\mu_{2}}{max}\hskip 14.22636pt\leavevmode\nobreak$ $e_{1}^{t}\alpha_{2}-\frac{1}{2}(\alpha_{2}^{t}G-\mu_{2}^{t}F)(EE^{t})^{-1}(G^{t}\alpha_{2}-F^{t}\mu_{2})$
$+(\epsilon-1)e_{u}^{t}\mu_{2}$
s.t $0\leq\mu_{2}\leq c_{u},$
$0\leq\alpha_{2}\leq c_{2},$
here, $G=[X_{1},e_{1}]$, $E=[X_{2},e_{2}]$, $F=[U,e_{u}]$ and $\alpha_{1}$, $\alpha_{2}$, $\mu_{1}$, $\mu_{2}$ are Lagrange multipliers of appropriate dimensions. The solution of (II-A) and (II-A) yields the separating hyperplane,
[…]
III Proposed Intuitionistic Fuzzy Universum Twin Support Vector Machine for Imbalanced Data (IFUTSVM-ID)
In this section, we provide a detailed mathematical formulation of the proposed IFUTSVM-ID model tailored for linear and nonlinear cases. The proposed IF-edRVFL incorporates an intuitionistic fuzzy scheme, enabling the development of multiple diverse and robust base models. In case of class imbalance, the majority class holds a significant amount of information regarding the data distribution. When dealing with a majority-class hyperplane, we increase the number of data points in the universum data to proper information about the data. In contrast to UTSVM, which does not address the issue of class imbalance and considers the universum as not belonging to either of the two classes. Constraints on the universum samples are included such that their sizes are equivalent to the difference between the number of samples in the classes. To construct the hyperplane for the majority class, a constraint is applied to ensure that the distance of universum data points from the majority (negative) class must be $(1-\epsilon)$. In addition to giving prior information of data, the majority class hyperplane is not positioned too close to the data point of the minority (positive) class. Consequently, this scheme reduces the inclination of the classifier towards the majority class, leading to improved generalization performance.
III-A Linear IFUTSVM-ID
$\underset{w_{1},b_{1},\xi_{1},\lambda_{1}}{min}\hskip 14.22636pt\leavevmode\nobreak$ $\frac{1}{2}\|X_{1}w_{1}+e_{1}b_{1}\|^{2}+\frac{c_{3}}{2}(\|w_{1}\|^{2}+b_{1}^{2})$
$+c_{1}s_{2}^{t}\xi_{1}+c_{u}e_{g}^{t}\lambda_{1}$
s.t. $-\left(X_{2}^{*}w_{1}+e_{2}b_{1}\right)\geq e_{2}-\xi_{1},$
$\left(U^{*}w_{1}+e_{g}b_{1}\right)\geq(-1+\epsilon)e_{g}-\lambda_{1},$
$\hskip 17.07182pt\xi_{1}\geq 0,\lambda_{1}\geq 0,$ (12)
and
$\underset{w_{2},b_{2},\xi_{2},\lambda_{2}}{min}\hskip 14.22636pt\leavevmode\nobreak$ $\frac{1}{2}\|X_{2}w_{2}+e_{2}b_{2}\|^{2}+\frac{c_{4}}{2}(\|w_{2}\|^{2}+b_{2}^{2})$
$+c_{2}s_{1}^{t}\xi_{2}+c_{u}e_{d}^{t}\lambda_{2}$
s.t. $\left(X_{1}w_{2}+e_{1}b_{2}\right)\geq e_{1}-\xi_{2},$
$\left(Uw_{2}+e_{d}b_{2}\right)\geq(\epsilon-1)e_{d}-\lambda_{2},$
$\hskip 17.07182pt\xi_{2}\geq 0,\lambda_{2}\geq 0,$ (13)
The Lagrangian of equation (III-A) is written as:
$L(\Theta)$ $=\frac{1}{2}\|X_{1}w_{1}+e_{1}b_{1}\|^{2}+\frac{c_{3}}{2}(\|w_{1}\|^{2}+b_{1}^{2})+c_{1}s_{2}^{t}\xi_{1}$
$+c_{u}e_{g}^{t}\lambda_{1}-\alpha^{t}(-\left(X_{2}^{*}w_{1}+e_{2}b_{1}\right)-e_{2}+\xi_{1})$
By employing the K.K.T. conditions to equation (III-A), we obtain
$X_{1}^{t}(X_{1}w_{1}+e_{1}b_{1})+c_{3}w_{1}+X_{2}^{*t}\alpha-U^{*t}\beta=0,$ (15)
$e_{1}^{t}(X_{1}w_{1}+e_{1}b_{1})+c_{3}b_{1}+e_{2}^{t}\alpha-e_{g}^{t}\beta=0,$ (16)
$c_{1}s_{2}-\alpha-\mu=0,$ (17)
$c_{u}e_{g}-\beta-\psi=0,$ (18)
$\alpha^{t}(-\left(X_{2}^{*}w_{1}+e_{2}b_{1}\right)-e_{2}+\xi_{1})=0,$ (19)
$\mu^{t}\xi_{1}=0,$ (21)
$\gamma^{t}\lambda_{1}=0.$ (22)
Since $\mu>0$ and $\gamma>0,$ (17) and (18) reduce to,
$0\leq\alpha\leq c_{1}s_{2},$
$0\leq\beta\leq c_{u}e_{g}.$
Rewriting (15) and (16), we obtain
$\left(E^{t}E+c_{3}I\right)\binom{w_{1}}{b_{1}}+F^{*t}\alpha-G^{*t}\beta$
[…]
With the help of equation (III-A) and the K.K.T conditions (15) - (22), the dual form of equation (III-A) can be obtained as follows:
$\underset{\alpha,\beta}{max}\hskip 14.22636pt\leavevmode\nobreak$ $-\frac{1}{2}\left(\alpha^{t}F^{*}-\beta^{t}G^{*}\right)(E^{t}E+c_{3}I)^{-1}(F^{*t}\alpha-G^{*t}\beta)$
$+(\epsilon-1)e_{g}^{t}\beta+e_{1}^{t}\alpha$
s.t. $0\leq\alpha\leq c_{1}s_{2},$
$0\leq\beta\leq c_{u}e_{g},$ (25)
and
$\underset{\eta,\theta}{max}\hskip 14.22636pt\leavevmode\nobreak$ $-\frac{1}{2}\left(\eta^{t}E+\theta^{t}G\right)(F^{t}F+c_{4}I)^{-1}(E^{t}\eta+G^{t}\theta)$
$+(1-\epsilon)e_{g}^{t}\theta+e_{1}^{t}\eta$
s.t. $0\leq\eta\leq c_{2}s_{1},$
$0\leq\theta\leq c_{u}e_{d},$ (26)
After solving (III-A) and (III-A), we can acquire the separating hyperplanes as follows
[…]
III-B Non-linear IFUTSVM-ID
$\underset{w_{1},b_{1},\xi_{1},\lambda_{1}}{min}\hskip 14.22636pt\leavevmode\nobreak$ $\frac{1}{2}\|\mathscr{K}(X_{1},D^{t})w_{1}+e_{1}b_{1}\|^{2}+\frac{c_{3}}{2}(\|w_{1}^{2}\|+b_{1}^{2})$
$+c_{1}s_{2}^{t}\xi_{1}+c_{u}e_{g}^{t}\lambda_{1}$
$\hskip 17.07182pt\xi_{1}\geq 0,\lambda_{1}\geq 0,$ (30)
and
$\underset{w_{2},b_{2},\xi_{2},\lambda_{2}}{min}\hskip 14.22636pt\leavevmode\nobreak$ $\frac{1}{2}\|\mathscr{K}(X_{2},D^{t})w_{2}+e_{2}b_{2}\|^{2}+\frac{c_{4}}{2}(\|w_{2}^{2}\|+b_{2}^{2})$
$+c_{2}s_{1}^{t}\xi_{2}+c_{u}e_{d}^{t}\lambda_{2}$
s.t. $\left(\mathscr{K}(X_{1},D^{t})w_{2}+e_{1}b_{2}\right)\geq e_{1}-\xi_{2},$
$\hskip 17.07182pt\xi_{2}\geq 0,\lambda_{2}\geq 0,$ (31)
The Wolfe duals of the problems of (III-B) and (III-B) can be obtained as:
$\underset{\alpha,\beta}{max}\hskip 14.22636pt\leavevmode\nobreak$ $-\frac{1}{2}\left(\alpha^{t}F^{*}-\beta^{t}G^{*}\right)(E^{t}E+c_{3}I)^{-1}(F^{*t}\alpha-G^{*t}\beta)$
$+(\epsilon-1)e_{g}^{t}\beta+e_{1}^{t}\alpha$
s.t. $0\leq\alpha\leq c_{1}s_{2},$
$0\leq\beta\leq c_{u}e_{g},$ (32)
and
$\underset{\eta,\theta}{max}\hskip 14.22636pt\leavevmode\nobreak$ $-\frac{1}{2}\left(\eta^{t}E+\theta^{t}G\right)(F^{t}F+c_{4}I)^{-1}(E^{t}\eta+G^{t}\theta)$
$+(1-\epsilon)e_{g}^{t}\theta+e_{1}^{t}\eta$
s.t. $0\leq\eta\leq c_{2}s_{1},$
$0\leq\theta\leq c_{u}e_{d},$ (33)
After solving (III-B) and (III-B), the separating hyperplanes can be attained as,
[…]
The algorithm and time complexity analysis of the proposed IFUTSVM-ID model are detailed in Section S.II of the supplementary material.
IV Experimental results
To test the efficiency of proposed IFUTSVM-ID model, we compare their performance against baseline models, including UTSVM [16], RUTSVM-CIL [31], IF-USVM [32], IFW-LSTSVM [33], and CGFTSVM-ID [34]. We utilize the publicly available KEEL [35] benchmark datasets under different noise levels. Furthermore, to assess the effectiveness of the proposed IFUTSVM-ID model, we utilized it for the diagnosis of Alzheimer’s Disease (AD). AD dataset is accessible through the Alzheimer’s Disease Neuroimaging Initiative (ADNI) (adni.loni.usc.edu). In the Supplementary Material, we discuss the experimental setup and the experimental results on KEEL datasets with label noise.
Table I: Average accuracy (ACC) and average rank of the proposed IFUTSVM-ID model with the baseline models based on KEEL datasets.
Metric $\downarrow$   Model $\rightarrow$ UTSVM [16] RUTSVM-CIL [31] IF-USVM [32] IFW-LSTSVM [33] CGFTSVM-ID [34] IFUTSVM-ID
Average ACC $83.52$ $81.15$ $80.61$ $79.44$ $86.19$ $87.70$
Average Rank $3.71$ $4.24$ $4.47$ $4.33$ $2.76$ $1.54$
IV-A Evaluation on KEEL datasets
In this subsection, we present an intricate analysis involving a comparison of the proposed IFUTSVM-ID model against baseline UTSVM, RUTSVM-CIL, IF-USVM, IFW-LSTSVM, and CGFTSVM-ID models. This comparison is carried out across $46$ KEEL benchmark datasets spanning diverse domains. Table S.II presents the classification ACC, along with the optimal parameters of the proposed IFUTSVM-ID model compared to the existing models in the Supplementary material. Table S.II shows the average ACC and the average Rank of the proposed IFUTSVM-ID model compared to the existing models. The average ACC of the existing UTSVM, RUTSVM-CIL, IF-USVM, IFW-LSTSVM, and CGFTSVM-ID models are $83.52\%$, $81.15\%$, $80.61\%$, $79.44\%$, and $86.19\%$ respectively, whereas the average ACC of the proposed IFUTSVM-ID model is $87.70\%$. The proposed IFUTSVM-ID secured the top position in terms of average ACC. This observation strongly emphasizes the significant superiority of the proposed models over baseline models. Depending solely on average ACC as a single metric could pose problems, as outstanding performance on specific datasets may obscure insufficient performance on others. To address this concern, it becomes crucial to rank each model individually with respect to each dataset, enabling a thorough evaluation of their respective capabilities. In the ranking scheme [36], the model with the lowest performance on a dataset receives a higher rank, whereas the model achieving the best performance is assigned a lower rank. In the evaluation of $p$ models across $N$ datasets, the rank of the $j^{th}$ model on the $i^{th}$ dataset can be denoted as $R_{j}^{i}$. Then, $R_{j}=\frac{1}{N}\sum_{i=1}^{N}R_{j}^{i}$ determined the average rank of the model. The average rank of the proposed IFUTSVM-ID with the existing UTSVM, RUTSVM-CIL, IF-USVM, IFW-LSTSVM, and CGFTSVM-ID models are $3.71$, $4.24$, $4.47$, $4.33$, $2.76$ and $1.54$, respectively. The proposed IFUTSVM-ID model attained the lowest average rank among all the models. Since a lower rank indicates a better-performing model, the proposed IFUTSVM-ID emerged as the top-performing model. We now proceed to conduct statistical tests to determine the significance of the results. Firstly, we employ the Friedman test [36] to determine whether there are significant differences among the models. Under the null hypothesis, it is assumed that the average rank of the models is equivalent, indicating equal performance. The Friedman statistic is given by: $\chi_{F}^{2}=\frac{12N}{p(p+1)}\left[\sum_{i=1}R_{j}^{2}-\frac{p(p+1)^{2}}{4}\right]$. However, the Friedman statistic tends to be overly cautious. To address this, the better statistic [37] is given as follows: $F_{F}=\frac{(N-1)\chi_{F}^{2}}{N(p-1)-\chi_{F}^{2}}$, where $F_{F}$ follows an $F$-distribution with d.o.f $((p-1),(p-1)(N-1))$. For $p=6$ and $N=46$, we get $\chi_{F}^{2}=91.48$ and $F_{F}=29.72$. From the $F$-distribution table $F_{F}(5,225)=2.2542$ at a $5\%$ level of significance. The null hypothesis rejected as $F_{F}>2.2542$. Therefore, significant disparities exist among the models. Subsequently, we utilize the Nemenyi post hoc test [36] to scrutinize the pairwise distinctions between the models. The critical difference $(C.D.)$ is computed at a significance level of $5\%$ as $C.D.=q_{\alpha}\sqrt{\frac{p(p+1)}{6N}}$, where $q_{\alpha}$ denotes the critical value obtained from the distribution table for the two-tailed Nemenyi test. Referring to the statistical $F$-distribution table, at a significance level of $5\%$, we find $q_{\alpha}=2.850$, which leads to a critical difference of $C.D.=1.11$. The average differences in rank between the proposed IFUTSVM-ID model and the baseline UTSVM, RUTSVM-CIL, IF-USVM, IFW-LSTSVM, and CGFTSVM-ID models are $2.17$, $2.70$, $2.93$, $2.79$, and $2.76$, respectively. The Nemenyi post hoc test reveals significant differences between the proposed IFUTSVM-ID model and the baseline models. Considering all these findings, we infer that the proposed IFUTSVM-ID model demonstrates superior performance when compared to the baseline models.
IV-B Evaluation on ADNI dataset
Table II: Performance comparison of the proposed IFUTSVM-ID along with the baseline models based on classification ACC on AD and BC datasets.
Average ACC $76.12$ $66.92$ $69.48$ $56$ $76.81$ $77.57$
The ACC values for the proposed IFUTSVM-ID model along with the existing models for AD diagnosis, are presented in Table II. With an average ACC of $77.57\%$, the proposed IFUTSVM-ID emerges as the top classifier in our analysis. The average ACC of the baseline UTSVM, RUTSVM-CIL, IF-USVM, IFW-LSTSVM, and CGFTSVM-ID models are $76.12$, $66.92$, $69.48$, $56$, and $76.81$, respectively. Compared to the second-top model, CGFTSVM-ID, the proposed IFUTSVM-ID model exhibits an average ACC that is approximately $0.76\%$ higher. The proposed IFUTSVM-ID achieves the highest ACC of $89.87\%$ for the CN vs AD case. In the CN vs MCI case, the proposed IFUTSVM-ID emerges as the top performer, with average accuracies of $70.85\%$. Therefore, the proposed IFUTSVM-ID consistently demonstrates superior performance by achieving high ACC across various cases, establishing its prominence among the models. Furthermore, to illustrate the comparison between the proposed IFUTSVM-ID model and the baseline UTSVM, RUTSVM-CIL, IF-USVM, IFW-LSTSVM, and CGFTSVM-ID models in terms of specificity, sensitivity, and precision, we plotted the bar graphs shown in Figure 2. The performance of the proposed IFUTSVM-ID model is almost comparable to that of the baseline models. The overall findings emphasize the effectiveness of the proposed models in discerning various cognitive states.
V Conclusion
In this paper, we proposed the intuitionistic fuzzy universum twin support vector machine for imbalanced data (IFUTSVM-ID) to enhance their robustness against noise and outliers present in the dataset. Unlike UTSVM, which employs a uniform weighting approach to generate the optimal classifier, the proposed IFUTSVM-ID model employs an intuitionistic fuzzy membership scheme to compute the final parameters. As a result, the proposed IFUTSVM-ID achieves superior generalization performance by incorporating prior information into the training process. The approach integrating undersampling with oversampling using universum data has been found to be beneficial in classifying imbalanced datasets. One of the key advantages of incorporating a regularization term into the primal optimization formulation of IFUTSVM-ID is its adherence to the SRM principle. This approach bypasses the need for matrix inversions, thereby streamlining computational processes. Furthermore, it effectively mitigates overfitting concerns, enhancing the model’s generalization capabilities. We evaluate the proposed IFUTSVM-ID model on $46$ KEEL datasets to experimentally validate the performance. Statistical evaluations, including metrics such as accuracy (ACC), ranking scheme analysis, Win-tie-loss sign test, Friedman test, and Nemenyi post hoc test, conclusively demonstrate that the proposed IFUTSVM-ID model exhibits superior performance compared to other classifiers. Also, we employed the proposed IFUTSVM-ID model for the diagnosis of Alzheimer’s disease datasets. Upon analysis, it has been ascertained that the proposed IFUTSVM-ID model demonstrates the highest level of classification performance. In future, there is a possibility of examining the selection of universum data as it has an impact on the performance of the model. Further, to enhance the performance of our IFUTSVM-ID model, various undersampling and oversampling techniques can be employed.
Acknowledgment
Supplementary Material
S.I Intuitionistic Fuzzy Membership Scheme